OpenStudy (anonymous):

Using the given zero, find all other zeros of f(x). -2i is a zero of f(x) = x4 - 21x2 - 100 @ganeshie8

OpenStudy (solomonzelman):

let x^2=a

OpenStudy (anonymous):

is is -2i is a zero of f(x) = x^4 - 21x^2 - 100

OpenStudy (solomonzelman):

\(\Large\color{blue}{ \bf hint.~~~~a^2-21a-100=(a -25)(a+4)}\)

OpenStudy (anonymous):

2i, 5, -5

OpenStudy (anonymous):

@SolomonZelman right?

OpenStudy (anonymous):

@ganeshie8 help please!!

ganeshie8 (ganeshie8):

Correct !

OpenStudy (solomonzelman):

\(\large\color{black}{ \bf x^2=25,~-4}\) THUS, \(\large\color{black}{ \bf x=±5,±2i}\) just -2i left out, but for other than that, you did a very good job, I actually like that :)